Back to the NIKA2 main page

Daily Reports

Because of the pandemic, no observers can come to the telescope. This pool is entirely in remote. First Pool week: Nicolas Ponthieu, Alexandre Beelen, Juan Macias, Laurence Perotto, Florian Kéruzoré, Stavroula Katsioli, Hélène Roussel, Florian Ruppin. A new sweep was performed.

Oct 27th, Tuesday

00h00 - 03h00; Florian K.: Together Ángel performed a pointing session from 23:15 to 01:15 (UT) - 61 raw sources. Went back to 199-16 for a repetition of observe_ACTJ0240.

03h00 - 09h00; Stavroula and Hélène: Opacities between ~ 0.3 and 0.5. Observed IC443G for project 037-20. Toward the end of the shift, opacities are quickly rising and conditions degrade. Note in TAPAS for 037-20 that the latest scans may be unusable and should be checked. Desired total integration time not yet achieved. A calib_1scan on Uranus and skydip were carried out.

09h00 - 10h30; Matthieu: Stopped at 09h00 because of high opacity. Maintenance works since then.

16h00 - 21h00; Ángel: Resuming from Maintenance works. High opacity (0.7<tau<1, strong instability). Switching to EMIR at ~18h for stability monitoring. Backup observations with EMIR (project 073-20) with slow opacity decreasing. Resuming with NIKA2 at ~21h00.

21h00 - 24h00; Miren, Jean-François & Juan: Pretty bad weather. Opacity at about 0.6. We observed 029-20 following Bilal instructions.

Oct 28th, Wednesday

00h00 - 02h00; Miren, Jean-François & Juan: Pretty bad weather. Opacity at about 0.6. We observed one more hour for 029-20 following Bilal instructions until source too low. Stopped since ~02h because of high opacity.

03h00 - 09h00; Stavroula and Hélène: Weather degrades again, with opacities well above 0.4 and unstable atmosphere.

09h00 - 15h00; Matthieu: On-hold. No observations. Bad weather.

15h00 - 21h00; F-Xavier Désert: opacity at 0.6 (but decreasing). Backup observations with EMIR (Mónica: project 153-19). We start observing with NIKA2 at 17 UT. Lots of anomalous refraction, going on and off. Moving to different pointing sources to get a decent focus.

21h00; Jean-François, Miren, Juan: relative good weather opacity around 0.3->0.2. Started observing MOO (Ruppin et al) and XXL clusters (Ricci et al). Good beammap on Uranus at around El.=0.60 and opacity 0.3. We observe since yesterday that box 13 is fully flagged by nk_rta, this should be investigated.

Oct 29th, Thursday

00h00 - 03h00; Jean-François, Miren, Juan: We observe ACTJ0240 for LPSZ project 199-16

03h00 - 09h00; Isabelle and Karine (assisted by Juan): We have been observing for our project 007-20 the following sources: L1506 and G150. Everything was fine, with good weather conditions :) We have commented on the scans in TAPAS indicating which ones should have been taken on nikas-20 account (pointing, focus, and skydip). Hopefully, we will not do the same mistake tomorrow.

09h00 - 03h00; Guilaine: starting the shift with opacity below 0.2; COSMOS field (192-16) and id34 (075-20) observed. Opacity is continuously increasing above 0.3 with a lot of instabilities. Numerous pointings and focus for about 1h30 with very large FWHMs and no good focus can be found.

15h00 - 21h00; FXD: opacity at 0.35 and rising. Lots of turbulence. Stay tuned. Handover to EMIR for a while. Backup observations with EMIR (Mónica: project 153-19). Resuming with Marco at 20h. Start 077-20 project on MOOJ0105.

21h00 - 24h00; Laurence and Jean-François: opacity stabilizing around 0.23. We observed 039_20, after a calib_1scan on Uranus, right after a focus-pointing sequence

Oct 30th, Friday

00h00 - 03h00; Laurence and Jean-François: opacity continued being stable around 0.23, while TOI showed large variations (hint of atmospheric instabilities ?); pointings yielded nice profiles and no indication of anomalous refraction (no shift between the scan in one direction and the scan backward). We observed 199-16 GT program. And finished the shift with a skydip and a calib-1scan on Uranus. We returned the antenna to the next team after pointing & focus.

03h00 - 09h00; Isabelle and Karine: starting shift with opacity ~0.2. We observed targets G150 and L1506 (007-20). After that, we mapped the GRB201024 region as requested in the project D04-20, with a mean tau~0.3 (and increasing). Then, we observed TMC1-C (above El. 24 deg.) of project 006-20. Apart from standard pointing and focus calibrations, two calib_1scan on Uranus and one skydip were launched. Finishing with opacity ~0.35.

09h00 - 15h00; Nicolas: 10h30: After small set up issues, because the opacity is still above 0.3 and increasing, we have switched to project 061-20 whose tau constraints are compatible with the current weather. Thanks to Angel's help in setting up scripts and paths, after a focus and pointing sequence, I'm completing right now the 3rd round of requested observations on JINGLE10. 12h02: Going to 1226+023 to have a pointing-focus source close to JINGLE184. The pointing gives rather large corrections but the beams are nice, hence no need to refocus. I produced the script observer_JINGLE184.pako from observer_JINGLE10.pako as it was missing and just launched it. 13h10: the telescope stopped for an unknown reason after the 1st loop on JINGLE184 (i.e. scan stuck). Angel and Sergio are working to solve the issue. Meanwhile, tau has kept increasing, so we decide to switch to EMIR.

14h30; Angel: the telescope responds again. Monitoring NGC7027 using EMIR: pwv~8 mm. Waiting for field rising of the backup EMIR project 153-19 (Mapping the North America Nebulae).

15h00 - 24h00; Mónica: Observing for project 153-19 until 18:30. Then, changed to the backup EMIR (3mm) project 011-20 (observation of starless-cores). OTF-maps on 1333-C1-1, B213-C5-1, B213-C6-1, B213-C7-1, and B213-C10-1. 0.45 < Tau_225Ghz < 0.55

On-hold to observe with NIKA2:

Oct 31st, Saturday

00h00 - 03h00; Wonju: Continuing 011-20. OTF-maps on B213-C10-10 and B213-C16. 0.47 < Tau_225Ghz < 1.48

03h00 - 05h15; Angel: Continuing 011-20. OTF-maps on B213-C10-10 and B213-C16. 0.65 < Tau_225Ghz < 0.95. Project 011-19 finished. No additional sources of backup EMIR (3 mm only) projects above the horizon until ~13h.

05h30 - 12h45; From taumeter, 0.7 < Tau_225Ghz < 0.87.

12:45 - 18h00; Monica: Start observing for backup EMIR project 153-19. Target fields at El. ~22 deg. when the instrument setup was launched.

18h00 - 21:40; Miguel: Observing source 919B of backup EMIR project 073-20. Tau_225GHz < 0.7 with spikes of +/-0.15.

21h40 - 24h00; F-Xavier: Resuming observations with NIKA2 at 21h30. After tuning the resonances, go to project 029-20 as opacity is above 0.6 but rather stable conditions.

On-hold to observe with NIKA2:

Nov 1st, Sunday

00h00 - 03h00; F-Xavier: Go to project 006-20 after calibration on Uranus. One hour on each, TMC1-CP and TMC1-C.

03h00 - 09h00; Angel: Continuing in the project 006-20. One hour on each source: TMC1-CP and TMC1-C. Quite good beam sizes from pointing on Uranus and 0439+60: <FWHM>_1mm,2mm = 11.6,17.7 arcsec. Pointing corrections/beam widths from PIIC-QL (updated DAFS) and the IDL pipeline were consistent within ~0.2 arcsec. At 07h02 we went to the project D04-20. Mean tau_225GHz before sunrise: 0.45 and going down slowly. Stable since 04h.

09h00 - 15h00; Longji: Starting shift with opacity ~0.4, which maintained with slight variations until 12h when fluctuated between 0.45 and 0.65. Despite this, beam sizes (2mm) measured on pointing scans are all below 18 arcsec. Targets observed JINGLE184 and JINGLE16 (061-20).

15h00 - 21:00; F-Xavier: Tau is decreasing from 0.6 to 0.4 but the beams are very unstable from 13 to 20 arcsec at 1mm and 17.5 to 20 at 1mm. We perform a series of calib_1scan on standard secondary calibrators before moving to 077-20 (because 029-20 is at too high elevation). 2 repeats on MOOJ2206 are performed (2 times 16 scans).

21h00 - 24h00; Florian K: Started by finishing MOOJ0105 for 077-20. Ran secondary calibrators and a beammap on Uranus (20201101s297).

Nov 2nd, Monday

00h00 - 03h00; Florian K: Two repetitions were obtained on IC443G for project 037-20. Opacity maintained about 0.3 during the shift.

03h00 - 09h00; Hélène and Matt: Opacities between ~ 0.3 and 0.4 from 3am to ~ 8am (not sure of the exact time), decreasing below 0.3 (above 0.25) at the end of the shift. Performed a calib_1scan on Uranus, 12 scans on IC443G (project 037-20) that is now roughly half-completed, 12 scans on GRB201024 (project D04-20, third epoch), and 9 scans on NGC2841 (project 160-16). The beams seemed good and the focus stable throughout the shift.

09h00 - 15h00; Longji: 3 additional scans on NGC2841 (project 160-16) completed. Started observing target JINGLE0 (061-20) with tau~0.25. Some instabilities appeared starting s112. Pointing corrections from PIIC-QL and RTA may differ even by 2 arcsec, focus corrections up to 0.2 mm.

15h00 - 21h00; Marco and Antonio: The shift started with increasing tau values (above 0.4) and large instabilities. Systematic monitoring of pointing sources, waiting for tau decreasing, reliable pointing/focus, and decent beam sizes. Impossible to try sky fields of 029-20 (tau < 0.8) before 18h because were at El. > 67 deg. After weather stabilization, the science observations were resumed at 18:45 (UT) doing one repetition on MOOJ2206 (077-20) and started the second before leaving for Florian K.

21h00 - 24h00; Florian K: Two additional repetitions on MOOJ2206 (077-20) were performed. I also made a beammap on Uranus.

Nov 3rd, Tuesday

00h00 - 03h00; Florian K: Coming back to project 077-20: three repetitions of observe_MOOJ0319 script were launched.

03h00 - 10h30; Ioannis and Angel: we continued observations of MOOJ0319 started by Florian, adding two repetitions. After that, we visited GRB201024 (D04-20) and a new epoch was added. Then a skydip (s99) was obtained. After focus+pointing (beam sizes: 11.4, 17.8 arcsec) we jump to IC443G (project 037-20; one repetition). Finishing the shift with a calib_1scan on 0851+202. We tried the same with 1226+023 before diurnal tau increasing, but the scan got stuck without immediate recovering. Telescope stopped at 10:05 for weekly maintenance.

DailyReportsNika2Pool271020 (last edited 2020-11-05 13:52:08 by NikaBolometer)